OpenStudy (anonymous):

Use the quadratic formula to solve x2 + 8x = –2.

OpenStudy (anonymous):

Solve for x over the real numbers: \[x^2+8 x = -2\] Take one half of the coefficient of x and square it, then add it to both sides. Add 16 to both sides: \[x^2+8 x+16 = 14\] Factor the left hand side. Write the left hand side as a square: \[(x+4)^2 = 14\] Eliminate the exponent on the left hand side. Take the square root of both sides: \[x+4 = sqrt(14) \] Look at the first equation: Solve for x.

OpenStudy (espex):

@fernmndez that is completing the square, the instructions state to use the quadratic formula. \[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

@eSpeX :( Sorry. I might've had a minor stroke. Not sure how I missed that..

OpenStudy (goformit100):